OpenStudy (anonymous):

If the distance between two objects is increased to four times the original distance, how will that increase change the force of attraction between the objects?

OpenStudy (shamim):

Gravitational attraction force is inversely proportional to the r^2

OpenStudy (shamim):

Here r is the distance between 2 objects

OpenStudy (shamim):

If r becomes 4r, the (4r)^2=16r^2

OpenStudy (shamim):

So gravitational attraction force will b decrease by 16 times

OpenStudy (shamim):

Response plz!

OpenStudy (michele_laino):

Hint: before to increase the distance, the force is: \[\Large F = K\frac{{{M_1}{M_2}}}{{{d^2}}}\] after that the distance has been increased, we can write: \[\Large {F_1} = K\frac{{{M_1}{M_2}}}{{{{\left( {4d} \right)}^2}}} = K\frac{{{M_1}{M_2}}}{{16{d^2}}} = \frac{F}{{16}}\]

OpenStudy (michele_laino):

where, as usually, K=6.67*10^(-11) N*m^2*Kg^2